OpenStudy (bananas1234):

evaluate the function g(t) = 2/3t, when t = 30

OpenStudy (mathmale):

This problem, g(t) = 2/3t, may seem deceptively simple. First of all, you have to determine whether the Original problem was g(t) = (2/3) t or g(t) = 2 / (3t). I suspect it's the former and not the latter that is correct. Choose whichever form you believe is correct, and then subst. t=30.

OpenStudy (jackthegreatest):

ahh good point

OpenStudy (bananas1234):

is 20 correct? @mathmale